# Generalization vs Specialization under Concept Shift

### v1

@@ Abstract
Machine learning models are often brittle under distribution shift, i.e., when data distributions at test time differ from those during training. Understanding this failure mode is central to identifying and mitigating safety risks of mass adoption of machine learning. Here we analyze ridge regression under concept shift—a form of distribution shift in which the input-label relationship changes at test time. We derive an exact expression for prediction risk in the high-dimensional limit. Our results reveal nontrivial effects of concept shift on generalization performance, depending on the properties of robust and nonrobust features of the input. We show that test performance can exhibit a nonmonotonic data dependence, even when double descent is absent. Finally, our experiments on MNIST and FashionMNIST suggest that this intriguing behavior is present also in classification problems.
@@ 1 Distribution shift
It is unsurprising that a model trained on one distribution does not perform well when applied to data from a different distribution. Yet, this out-of-distribution setting is relevant to many practical applications from scientific research [1, 2] to medicine and healthcare [3, 4, 5]. A quantitative understanding of out-of-distribution generalization is key to developing safe and robust machine learning techniques.
A model that generalizes to arbitrary distribution shifts of course does not exist. The scope of a model, however, needs not be limited to the training data distribution. A question then arises as to how much a model’s scope extends beyond its training distribution.
@@
Here we consider concept shift—a relatively less-studied setting, in which the input distribution is preserved but the input-label relationship becomes different at test time [11, 12]. We formulate a minimal model which enables continuous modulation of the input-label function, based on high dimensional ridge regression—a solvable setting that has helped develop intuitions for some of the most interesting phenomena of modern machine learning (e.g., Refs [13, 14, 15, 16, 17, 18, 19, 20, 21]). We derive an analytical expression for prediction risk under concept shift and demonstrate that its behavior can change from monotonically decreasing with data for the in-distribution case to monotonically increasing, and to nonmonotonic, depending on the degree of concept shift and the properties of the inputs. To isolate these effects from double descent phenomena [22, 23, 24], we focus on optimally tuned ridge regression which completely suppresses the risk divergence at the interpolation threshold and for which in-distribution prediction risk decreases monotonically with more data [13, 14]. Finally, we illustrate similar qualitative changes in test performance behavior in classification problems, using MNIST and FashionMNIST as examples. Our work contributes a new theoretical framework for analyzing concept shift that complements an extensive body of work on concept shift detection (see, e.g., Ref [25] for a recent review).
@@ 2 Regression setting
Data—The training data consists of $N$ iid input-response pairs $\{(x_{1},y_{1}),\dots,(x_{N},y_{N})\}$. The input $x\in\mathbb{R}^{P}$ is a vector of Gaussian features and the response $y\in\mathbb{R}$ is a noisy linear projection of $x$, i.e.,
@@
$\left[\begin{matrix}x\\ y\end{matrix}\right]\sim\mathcal{N}\left(\cdot,\left[\begin{matrix}\Sigma&\Sigma\beta\\ \beta^{\mkern-1.5mu\mathsf{T}}\Sigma&\beta^{\mkern-1.5mu\mathsf{T}}\Sigma\beta+\sigma_{\xi}^{2}\end{matrix}\right]\right)\quad\text{and}\quad\left[\begin{matrix}\tilde{x}\\ \tilde{y}\end{matrix}\right]\!\sim\!\mathcal{N}\left(\cdot,\left[\begin{matrix}\tilde{\Sigma}&\tilde{\Sigma}\tilde{\beta}\\ \tilde{\beta}^{\mkern-1.5mu\mathsf{T}}\tilde{\Sigma}&\tilde{\beta}^{\mkern-1.5mu\mathsf{T}}\tilde{\Sigma}\tilde{\beta}+\tilde{\sigma}_{\xi}^{2}\end{matrix}\right]\right),$ (2)
where in general $\Sigma\!\neq\!\tilde{\Sigma}$, $\beta\!\neq\!\tilde{\beta}$ and $\sigma_{\xi}^{2}\!\neq\!\tilde{\sigma}_{\xi}^{2}$.
Model—We consider ridge regression in which the predicted response to an input $x$ reads $\hat{y}(x;X,Y)\!=\!x\cdot\hat{\beta}_{\lambda}(X,Y)$, with the coefficient vector resulting from minimizing $L_{2}$-regularized mean square error,
$\hat{\beta}_{\lambda}(X,Y)\equiv\arg\min_{b\in\mathbb{R}^{P}}\tfrac{1}{N}\|Y-X^{\mkern-1.5mu\mathsf{T}}b\|^{2}+\lambda\|b\|^{2}=(XX^{\mkern-1.5mu\mathsf{T}}+\lambda NI_{P})^{-1}XY.$ (3)
@@
Risk—We measure generalization performance with prediction risk,
@@
$B(X)$ $\equiv\operatorname*{\mathbf{E}}\left[\left\|\operatorname*{\mathbf{E}}(\hat{y}(\tilde{x};X,Y)\mid X,\tilde{x})-\operatorname*{\mathbf{E}}(\tilde{y}\mid\tilde{x})\right\|^{2}\mid X\right]$ (5)
$V(X)$ $\equiv\operatorname*{\mathbf{E}}\left[\left\|\hat{y}(\tilde{x};X,Y)-\operatorname*{\mathbf{E}}(\hat{y}(\tilde{x};X,Y)\mid X,\tilde{x})\right\|^{2}\mid X\right].$ (6)
@@
$B(X)=\left(\tfrac{\Psi}{\Psi+\lambda I_{P}}\beta-\tilde{\beta}\right)^{\mkern-1.5mu\mathsf{T}}\tilde{\Sigma}\left(\tfrac{\Psi}{\Psi+\lambda I_{P}}\beta-\tilde{\beta}\right)\quad\text{and}\quad V(X)=\sigma_{\xi}^{2}\tfrac{1}{N}\operatorname{Tr}\left(\tilde{\Sigma}\tfrac{\Psi}{(\Psi+\lambda I_{P})^{2}}\right),$ (7)
@@
This intuitive picture breaks down under concept shift, $\tilde{\beta}\!\neq\!\beta$. Consider, for example, $\tilde{\beta}\!=\!0$ which indicates that none of the features predicts the response at test time. In this case, $R_{N=0}(X)\!=\!0$ and $R_{N\to\infty}(X)\!=\!\beta^{\mkern-1.5mu\mathsf{T}}\tilde{\Sigma}\beta\!>\!0$ (see also Fig 1); that is, even infinitely more data hurts test performance.
@@ 3 High dimensional limit
To better understand this counterintuitive phenomenon in the context of high dimensional learning, we focus on concept shift without covariate shift, i.e., $\tilde{\beta}\!\neq\!\beta$ and $\tilde{\Sigma}\!=\!\Sigma$, and take the thermodynamic limit $N,P\!\to\!\infty$ and $P/N\!\to\!\gamma\!\in\!(0,\infty)$. In this limit, prediction risk becomes deterministic $R(X)\!\to\!\mathcal{R}$ with the bias and variance contributions given by (see Appendix A for derivation; see also Ref [26]),
@@
where we define $\hat{G}_{\Sigma}(z)\!\equiv\!(m(z)\Sigma-zI_{P})^{-1}$, $m(z)\!\equiv\!(1+\gamma\nu(z))^{-1}$ and $\nu(z)\!\in\!\mathbb{C}^{+}$ is the solution of $\nu(z)\!=\!\frac{1}{P}\operatorname{Tr}[\Sigma\hat{G}_{\Sigma}(z)]$. We note that concept shift enters prediction risk only through the last two terms of the bias, Eq (8), whereas the variance, Eq (9), is completely unaffected.
@@ 3.1 Isotropic features
When $\Sigma\!=\!I_{P}$, the bias contribution reads (see Appendix B for the closed-form expression for $\nu(z)$)
@@
$\mathcal{B}=\|\beta\|^{2}\big{[}\lambda^{2}\nu^{\prime}(-\lambda)-2\lambda g(-z)(1-\kappa\cos\theta)+1+\kappa^{2}-2\kappa\cos\theta\big{]},$ (10)
where $g(z)\!\equiv\!(m(z)-z)^{-1}$ and we parametrize concept shift by the coefficient alignment $\cos\theta\!\equiv\!\beta\!\cdot\!\tilde{\beta}/\|\beta\|\|\tilde{\beta}\|$ and scaling factor $\kappa\!\equiv\!\|\tilde{\beta}\|/\|\beta\|$.
@@
Figure 1 depicts thermodynamic-limit prediction risk for isotropic features under concept shift. We focus on in-distribution optimal ridge regression which corresponds to setting $\lambda\!=\!\gamma/\textsc{snr}$, ensuring that double descent is absent (see, e.g., Refs [13, 14]). In Fig 1a, the coefficient vector becomes smaller at test time without changing direction, $\theta\!=\!0$ and $\kappa\!\leq\!1$. At $\kappa\!=\!1$, concept shift is absent and prediction risk monotonically decreases with more training data, as expected for optimally-tuned ridge regression. As $\kappa$ decreases and the features become less predictive of the response at test time, prediction risk starts to increase with training sample size. We observe a similar crossover in different settings. In Fig 1b, concept shift occurs only through misalignment, $\kappa\!=\!1$ and $\theta\!\geq\!0$. In Fig 1c, we consider a mixture of robust and nonrobust features with $\tilde{\beta}_{i}\!=\!\beta_{i}$ if feature $i$ is robust, otherwise $\tilde{\beta}_{i}\!=\!0$. More data hurts prediction risk in all of the above settings for adequately strong concept shift.
Figure 1: More data hurts performance when concept shift is strong. We depict the data dependence of asymptotic prediction risk for isotropic features, Eq (10), under three concept shift settings of varying degree, parametrized by coefficient alignment $\cos\theta\!=\!\beta\!\cdot\!\tilde{\beta}/\|\beta\|\|\tilde{\beta}\|$ and scaling factor $\kappa\!=\!\|\tilde{\beta}\|/\|\beta\|$ (see legend). a Shrinking coefficients: $\tilde{\beta}\!=\!\kappa\beta$. b Rotating coefficients: $\|\tilde{\beta}\|\!=\!\|\beta\|$ but $\theta$ varies. c Mixture of robust and nonrobust features: $\tilde{\beta}_{i}\!=\!\beta_{i}$ if feature $i$ is robust, otherwise $\tilde{\beta}_{i}\!=\!0$. This setting is parametrized by $q\!=\!\kappa^{2}\!=\!\cos^{2}\theta$. Here $\textsc{snr}\!=\!1$ and regularization is in-distribution optimal.

### v2

@@ Abstract
Machine learning models are often brittle under distribution shift, i.e., when data distributions at test time differ from those during training. Understanding this failure mode is central to identifying and mitigating safety risks of mass adoption of machine learning. Here we analyze ridge regression under concept shift—a form of distribution shift in which the input-label relationship changes at test time. We derive an exact expression for prediction risk in the thermodynamic limit. Our results reveal nontrivial effects of concept shift on generalization performance, including a phase transition between weak and strong concept shift regimes and nonmonotonic data dependence of test performance even when double descent is absent. Our theoretical results are in good agreement with experiments based on transformers pretrained to solve linear regression; under concept shift, too long context length can be detrimental to generalization performance of next token prediction. Finally, our experiments on MNIST and FashionMNIST suggest that this intriguing behavior is present also in classification problems.
@@ 1 Distribution shift
It is unsurprising that a model trained on one distribution does not perform well when applied to data from a different distribution. Yet, this out-of-distribution setting is relevant to many practical applications from scientific research [1, 2] to medicine and healthcare [3, 4, 5, 6]. A quantitative understanding of out-of-distribution generalization is key to developing safe and robust machine learning techniques. A model that generalizes to arbitrary distribution shifts of course does not exist. The generalization scope of a model, however, needs not be limited to the training data distribution. A question then arises as to how much a model’s scope extends beyond its training distribution.
Figure 1: Two flavors of distribution shift. Distribution shift describes the scenarios in which the joint input-label distribution of training data $\{(x_{1},y_{1}),(x_{2},y_{2}),\dots\}$ (filled circles) differs from that of test data $\{(\tilde{x}_{1},\tilde{y}_{1}),(\tilde{x}_{2},\tilde{y}_{2}),\dots\}$ (empty circles). Covariate shift (left) assumes a fixed input-label relationship but the input distribution differs at test time, i.e., $P_{Y|X}\!=\!P_{\smash{\tilde{Y}|\tilde{X}}}$ but $P_{X}\!\neq\!P_{\smash{\tilde{X}}}$. For linear regression, this condition means that the regression coefficient $\beta$ is unchanged. Concept shift (right) allows the input-label function to change at test time, which corresponds to a shift in regression coefficient $\beta\!\neq\!\tilde{\beta}$ in the linear regression setting. See §2.
@@
We consider concept shift—a relatively less-studied setting, in which the input-label relationship becomes different at test time [^1] Concept shift or concept drift is sometimes defined to be equivalent to distribution shift [12]. Here, we adopt a narrower definition in which concept shift describes only the change in the input-label relationship [13, 14].  [13, 14], see Fig 1. While many works have studied how to detect and mitigate concept shift [15, 16], characterizing how concept shift affects generalization behavior in neural networks has remained underexplored. In our work, we formulate a minimal model which enables continuous modulation of the input-label function, based on high dimensional ridge regression—a solvable setting that has helped develop intuitions for some of the most interesting phenomena of modern machine learning (see, e.g., Refs [17, 18, 19, 20, 21, 22, 23, 24, 25]). We derive an analytical expression for prediction risk under concept shift in the high-dimensional limit and demonstrate that its behavior can change from monotonically decreasing with data for the in-distribution case to monotonically increasing, and to nonmonotonic, depending on the degree of concept shift and the properties of the input distribution.
To differentiate these effects from double descent phenomena which can also cause nonmonotonic data dependence of prediction risk [26, 27], we focus on optimally tuned ridge regression which completely suppresses the risk divergence at the interpolation threshold and for which in-distribution prediction risk decreases monotonically with more data [17, 28, 18]. The nonmonotonic behavior we observe is also distinct from effects due to model misspecification [18]. While misspecification—a property of a model relative to the true data-generating process—can produce similar generalization behavior, distribution shift represents a fundamentally different setting in which the data-generating process changes at test time. Our work isolates and characterizes the effects of concept shift, a type of distribution shift, in correctly specified, optimally regularized models.
Our theoretical results in the thermodynamic limit agree well with experiments, based on transformers trained to solve finite-dimensional regression using in-context examples. We show that more in-context examples help improve model performance when concept shift is weak, but can lead to overspecialization for strong concept shift. Finally, we illustrate similar qualitative changes in generalization behavior in classification problems, using MNIST and FashionMNIST as examples. Our work contributes a new theoretical framework for analyzing concept shift that complements an extensive body of work on concept shift detection (see, e.g., Ref [29] for a recent review).
Our main contributions are:
  1. 1.
    We develop a theoretical framework, based on high-dimensional ridge regression, that isolates the effects of concept shift—an important yet often overlooked mode of distribution shift.
  2. 2.
    We demonstrate that more training data can harm generalization performance under adequately strong concept shift, characterizing a precise phase transition between weak and strong concept shift regimes across several settings.
  3. 3.
    We show that feature anisotropy creates qualitatively different patterns of risk nonmonotonicity depending on whether concept shift affects high or low-variance features.
  4. 4.
    Our experiments on transformers and simple classification tasks are in good qualitative agreement with our theoretical findings, hinting at universal behavior that generalizes beyond our relatively simple theoretical settings.
@@ 2 Regression setting
Data.
The training data consists of $N$ iid input-response pairs $\{(x_{1},y_{1}),\dots,(x_{N},y_{N})\}$. The input $x\in\mathbb{R}^{P}$ is a vector of Gaussian features and the response $y\in\mathbb{R}$ is a noisy linear projection of $x$, i.e.,
@@
Training data: $\left[\begin{matrix}x\\ y\end{matrix}\right]$ $\sim\mathcal{N}\left(\cdot,\left[\begin{matrix}\Sigma&\Sigma\beta\\ \beta^{\mkern-1.5mu\mathsf{T}}\Sigma&\beta^{\mkern-1.5mu\mathsf{T}}\Sigma\beta+\sigma_{\xi}^{2}\end{matrix}\right]\right)$ (2)
Test data: $\left[\begin{matrix}\tilde{x}\\ \tilde{y}\end{matrix}\right]$ $\sim\mathcal{N}\left(\cdot,\left[\begin{matrix}\tilde{\Sigma}&\tilde{\Sigma}\tilde{\beta}\\ \tilde{\beta}^{\mkern-1.5mu\mathsf{T}}\tilde{\Sigma}&\tilde{\beta}^{\mkern-1.5mu\mathsf{T}}\tilde{\Sigma}\tilde{\beta}+\tilde{\sigma}_{\xi}^{2}\end{matrix}\right]\right),$
where in general $\Sigma\!\neq\!\tilde{\Sigma}$, $\beta\!\neq\!\tilde{\beta}$ and $\sigma_{\xi}^{2}\!\neq\!\tilde{\sigma}_{\xi}^{2}$. Here we also define signal-to-noise ratio $\textsc{snr}\equiv\beta^{\mkern-1.5mu\mathsf{T}}\Sigma\beta/\sigma_{\xi}^{2}$
Model.
We consider ridge regression in which the predicted response to an input $x$ reads $\hat{y}(x;X,Y)\!=\!x\cdot\hat{\beta}_{\lambda}(X,Y)$, with the coefficient vector resulting from minimizing $L_{2}$-regularized mean square error,
$\hat{\beta}_{\lambda}(X,Y)\equiv\arg\min_{b\in\mathbb{R}^{P}}\frac{1}{N}\|Y-X^{\mkern-1.5mu\mathsf{T}}b\|^{2}+\lambda\|b\|^{2}=(XX^{\mkern-1.5mu\mathsf{T}}+\lambda NI_{P})^{-1}XY.$ (3)
@@
Risk.
We measure generalization performance with prediction risk,
@@
$B(X)$ $\equiv\operatorname*{\mathbf{E}}\left[\left\|\operatorname*{\mathbf{E}}(\hat{y}(\tilde{x};X,Y)\mid X,\tilde{x})-\operatorname*{\mathbf{E}}(\tilde{y}\mid\tilde{x})\right\|^{2}\mid X\right]$
$V(X)$ $\equiv\operatorname*{\mathbf{E}}\left[\left\|\hat{y}(\tilde{x};X,Y)-\operatorname*{\mathbf{E}}(\hat{y}(\tilde{x};X,Y)\mid X,\tilde{x})\right\|^{2}\mid X\right].$
@@
$B(X)=\left(\mbox{\small$\displaystyle\frac{\Psi}{\Psi+\lambda I_{P}}$}\beta-\tilde{\beta}\right)^{\mkern-1.5mu\mathsf{T}}\tilde{\Sigma}\left(\mbox{\small$\displaystyle\frac{\Psi}{\Psi+\lambda I_{P}}$}\beta-\tilde{\beta}\right)\quad\text{and}\quad V(X)=\sigma_{\xi}^{2}\mbox{\small$\displaystyle\frac{1}{N}$}\operatorname{Tr}\left(\tilde{\Sigma}\mbox{\small$\displaystyle\frac{\Psi}{(\Psi+\lambda I_{P})^{2}}$}\right),$ (5)
@@
This intuitive picture breaks down under concept shift, $\tilde{\beta}\!\neq\!\beta$. Consider, for example, $\tilde{\beta}\!=\!0$ which indicates that none of the features predicts the response at test time. In this case, $R_{N=0}(X)\!=\!0$ and $R_{N\to\infty}(X)\!=\!\beta^{\mkern-1.5mu\mathsf{T}}\tilde{\Sigma}\beta\!>\!0$ (see also Fig 2); that is, even infinitely more data decreases test performance in the presence of concept shift.
Figure 2: More data hurts performance when concept shift is strong. We depict the data dependence of asymptotic prediction risk for isotropic features, Eq (9), under three concept shift settings of varying degree, parametrized by coefficient alignment $\cos\theta\!=\!\beta\!\cdot\!\tilde{\beta}/\|\beta\|\|\tilde{\beta}\|$ and scaling factor $\kappa\!=\!\|\tilde{\beta}\|/\|\beta\|$ (see legend). a Shrinking coefficients: $\tilde{\beta}\!=\!\kappa\beta$. b Rotating coefficients: $\|\tilde{\beta}\|\!=\!\|\beta\|$ but $\theta$ varies. c Mixture of robust and nonrobust features: $\tilde{\beta}_{i}\!=\!\beta_{i}$ if feature $i$ is robust, otherwise $\tilde{\beta}_{i}\!=\!0$. This setting is parametrized by $q\!=\!\kappa^{2}\!=\!\cos^{2}\theta$. We set $\textsc{snr}\!=\!1$ in a-c, and the regularization is in-distribution optimal $\lambda\!=\!\gamma/\textsc{snr}$.
Figure 3: Concept shift induces a phase transition in data dependence behavior of generalization performance. We depict minimum prediction risk $\mathcal{R}_{\text{min}}$ of optimally tuned ridge regression ($\lambda\!=\!\gamma/\textsc{snr}$) as $\kappa$ and $\cos\theta$ vary, see Eq (10) for definitions. No concept shift corresponds to $\kappa\!=\!\cos\theta\!=\!1$ (star). The thick curve, $\kappa\cos\theta\!=\!1/2$, separates the weak and strong concept shift regimes. More training data improves generalization only when concept shift is weak, $\kappa\cos\theta\!>\!1/2$. Outside this region, any data hurts generalization.
@@ 3 High dimensional limit
To better understand this counterintuitive phenomenon in the context of high dimensional learning, we focus on concept shift without covariate shift, i.e., $\tilde{\beta}\!\neq\!\beta$ and $\tilde{\Sigma}\!=\!\Sigma$, and take the thermodynamic limit $N,P\!\to\!\infty$ and $P/N\!\to\!\gamma\!\in\!(0,\infty)$. In this limit, prediction risk becomes deterministic $R(X)\!\to\!\mathcal{R}$ with the bias and variance contributions given by (see Appendix for derivation; see also Ref [30]),
@@
where we define $\hat{G}_{\Sigma}(z)\!\equiv\!(m(z)\Sigma-zI_{P})^{-1}$, $m(z)\!\equiv\!(1+\gamma\nu(z))^{-1}$ and $\nu(z)$ is the unique solution of the self-consistent equation,
$\nu(z)=\mbox{\small$\displaystyle\frac{1}{P}$}\operatorname{Tr}[\Sigma\hat{G}_{\Sigma}(z)]\quad\text{with}\quad\nu(z)\in\mathbb{C}^{+}.$ (8)
We note that concept shift enters prediction risk only through the last two terms of the bias, Eq (6), whereas the variance, Eq (7), is completely unaffected.
@@ 3.1 Isotropic features
When $\Sigma\!=\!I_{P}$, the bias contribution to prediction risk reads (see Appendix for a closed-form expression for $\nu(z)$)
@@
$\mathcal{B}=\|\beta\|^{2}\Big{[}\lambda^{2}\nu^{\prime}(-\lambda)-\mbox{\small$\displaystyle\frac{2\lambda(1+\gamma\nu(-\lambda))}{1+\lambda(1+\gamma\nu(-\lambda))}$}(1-\kappa\cos\theta)+1-2\kappa\cos\theta+\kappa^{2}\Big{]},$ (9)
where we quantify concept shift via two parameters
[…]
Scaling factor: $\kappa$ $\equiv\|\tilde{\beta}\|/\|\beta\|.$
Figure 2 depicts thermodynamic-limit prediction risk for isotropic features under concept shift. We focus on in-distribution optimal ridge regression which corresponds to setting $\lambda\!=\!\gamma/\textsc{snr}$, ensuring that double descent is absent (see, e.g., Refs [17, 18]).
@@
In Fig 2a, we consider the effects of shrinking coefficients—the coefficient vector becomes smaller at test time without changing direction, $\theta\!=\!0$ and $\kappa\!\leq\!1$. When $\kappa\!=\!1$, concept shift is absent and prediction risk monotonically decreases with more training data, as expected for optimally-tuned ridge regression. As $\kappa$ decreases and the features become less predictive of the response at test time, prediction risk starts to increase with training sample size. This transition occurs at $\kappa\!=\!1/2$, at which $\mathcal{R}/\|\beta\|^{2}\!=\!1/4$.
In Fig 2b, we observe a similar crossover when the magnitude of the coefficient vector is fixed but its direction changes at test time, $\kappa\!=\!1$ and $\theta\!\geq\!0$. Examples of this concept shift setting include the case where some features have the opposite effects at test time, described by $\tilde{\beta}_{i}\!=\!-\beta_{i}$ for the affected coefficient. Here we see that more data hurts when $\cos\theta\!<\!1/2$ (or equivalently $\theta\!>\!\pi/3$).
In Fig 2c, we consider a mixture of robust and nonrobust features with $\tilde{\beta}_{i}\!=\!\beta_{i}$ if feature $i$ is robust, otherwise $\tilde{\beta}_{i}\!=\!0$. That is, the robust features have the same effects at test time whereas the nonrobut ones become uninformative of the response variables. In this case, we have $\kappa\!=\!\cos\theta\!=\!\sqrt{q}$ where $10\!\leq\!q\!\leq\!$ 1 denotes the fraction of robust features. We see again that adequately strong concept shift changes the data-dependent behavior of generalization properties from improving to worsening with more training data.
Figure 4: Concept shift generalization exhibits sample size nonmonotonicity for anisotropic features. We illustrate prediction risk for the two-scale model, Eq (11), with aspect ratio $s_{-}/s_{+}\!=\!0.1$, spectral weights $\rho_{+}\!=\!\rho_{-}\!=\!\nicefrac{{1}}{{2}}$ and signal fraction $\pi_{+}\!=\!\pi_{-}\!=\!\nicefrac{{1}}{{2}}$. Concept shift is parametrized by coefficient alignments $\cos\theta_{\pm}\!=\!\beta_{\pm}\!\cdot\!\tilde{\beta}_{\pm}/\|\beta_{\pm}\|\|\tilde{\beta}_{\pm}\|$ and scaling factors $\kappa_{\pm}\!=\!\|\tilde{\beta}_{\pm}\|/\|\beta_{\pm}\|$ where the subscripts indicate the variance $s_{\pm}$ of the affected features. We consider the settings in which concept shift affects either low or high-variance features and either alignment or scale; that is, we vary only one out of four parameters, $\theta_{\pm}$ and $\kappa_{\pm}$, at a time (see legend). a Shrinking coefficients of low-variance features. b Rotating coefficients of low-variance features. c and d Same as Panels a and b, but concept shift affects high-variance features via $\kappa_{+}$ and $\cos\theta_{+}$, respectively. Here $\textsc{snr}\!=\!1$ and regularization is in-distribution optimal.
Indeed we can make a more general statement about the transition between weak and strong concept shift. First, we observe that for in-distribution optimal ridge regression, $\lambda\!=\!\gamma/\textsc{snr}$, the thermodynamic-limit risk is always monotonic in $\gamma$. To determine whether more data improve generalization, we only need to compare the limits of no data and infinite data: $\mathcal{R}_{N=0}\!=\!\kappa^{2}\|\beta\|^{2}$ and $\mathcal{R}_{N\to\infty}\!=\!(1-2\kappa\cos\theta+\kappa^{2})\|\beta\|^{2}$. It follows immediately that two qualitatively distinct regimes exist. When $\kappa\cos\theta\!>\!1/2$, concept shift is weak and more data improves generalization. When $\kappa\cos\theta\!<\!1/2$, concept shift is strong and more data hurts, see Fig 3. When $\kappa\cos\theta\!=\!1/2$, prediction risk becomes completely independent of training sample size—i.e., $\mathcal{R}\!=\!\mathcal{R}_{N=0}\!=\!\mathcal{R}_{N\to\infty}\!=\!\kappa^{2}\|\beta\|^{2}$. Quite remarkably, this analysis does not depend on snr.333snr controls how prediction risk depends on training data size $N$ (Fig 2), but not the transition between weak and strong concept shift regimes (Fig 3).